OpenStudy (anonymous):

guys ........ x+y=2 and x-y=-2 how can i solve it

OpenStudy (hba):

what do u wanna find

OpenStudy (anonymous):

do it thru elimination!!

OpenStudy (anonymous):

what is elimination

OpenStudy (hba):

x+y=2 x-y=-2 2x=0 x=0

OpenStudy (anonymous):

HOW >??? i m confused

OpenStudy (hba):

0+y=2 y=2

OpenStudy (hba):

i am also confused

OpenStudy (anonymous):

add both the equations...y will be eliminated!!

mathslover (mathslover):

x+y=2-------------1) x-y=-2 ------------2) solving for x in 1st equation : x+y=2 x=2-y now put this value in 2nd equation x-y=-2 2-y-y=-2 2-2y=-2 -2y=-4 y=\(\frac{-4}{-2}\) y=2 x = 2-y = 2-2=0

OpenStudy (hba):

how did i do it

OpenStudy (anonymous):

didnt get you guys

mathslover (mathslover):

Try to study it ... from videos ..

OpenStudy (anonymous):

@hba and Hashir i am rather amazed at how fast memory lost and mental retardation can occur

OpenStudy (anonymous):

what is to understnd...add both the equations...2 x = 0

OpenStudy (anonymous):

@virtus : short term memory loss :D

OpenStudy (hba):

yeah right

OpenStudy (anonymous):

You can use a metrics if you don't understand elimination or substitution

OpenStudy (hba):

ghajni

OpenStudy (anonymous):

oh whats a matrix ... is it matrix .... the movie ... i love that movie

OpenStudy (anonymous):

:P

OpenStudy (ujjwal):

can't you add equatins @Hashir ?

OpenStudy (hba):

what is that

OpenStudy (anonymous):

@ujjwal : what equation ???

OpenStudy (anonymous):

equation ... i dont know what is an equation

OpenStudy (hba):